OpenStudy (anonymous):

indentify the vertex and the axis of symmery for the graph of y=3(x-3)^2+4

OpenStudy (anonymous):

= 3 (x^2 -6x +9) + 4 = 3x^2 - 18x + 27 + 4 = 3x^2 - 18x + 31 vertex = 18 / 6 = 3 axis of symmetry is x=3

OpenStudy (anonymous):

axis of symmetry is x = 3, the vertex is (3,4)

OpenStudy (anonymous):

yes sorry i didnt give the y cood of the vertex which is vertex = (3,4)